OpenStudy (anonymous):

Find the exact value of cos (5pi/12) by using a half-angle identity. 

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\frac{5\pi}{12}\] is half of \[\frac{5\pi}{6}\] use that one !!

OpenStudy (misty1212):

\[\cos(\frac{5\pi}{12})=\sqrt{\frac{1+\cos(\frac{5\pi}{6})}{2}}\]

OpenStudy (anonymous):

Hi misty 1212 :) Thanks for replying I know the formula is cos (a/2) = ±√(1 + cos a)/2 but I don't know what to do next

OpenStudy (misty1212):

compute the number is all that is left to do

OpenStudy (misty1212):

id you see the answer i wrote above? it was with \(\frac{a}{2}=\frac{5\pi}{12}\) and \(a=\frac{5\pi}{6}\)

OpenStudy (misty1212):

did i lose you somewheres?

OpenStudy (anonymous):

OK yea so would that be cos(5 pi/6) = -sqrt(3)/2?

OpenStudy (misty1212):

at this point what you need to do is find \[\cos(\frac{5\pi}{6})\] and plug it in above

OpenStudy (misty1212):

yes that is right

OpenStudy (misty1212):

\[\cos(\frac{5\pi}{12})=\sqrt{\frac{1+\cos(\frac{5\pi}{6})}{2}}=\sqrt{\frac{1-\frac{\sqrt3}{2}}{2}}\]

OpenStudy (misty1212):

you can simplify that compound fraction by multiplying inside the radical top and bottom by 2

OpenStudy (misty1212):

oh i should add that there is a plus or minus outside in the formula, but the answer is positive in this case because you are in quadrant 1

OpenStudy (anonymous):

\[2 \cos ^2\frac{ 5 \pi }{ 12 }=1+\cos \frac{ 5 \pi }{ 6 }=1+\cos \left( \pi-\frac{ \pi }{ 6 } \right)=1-\cos \frac{ \pi }{ 6 }=1-\frac{ \sqrt{3} }{ 2 }\] \[\cos \frac{ 5 \pi }{ 12 }=\sqrt{\frac{ 2-\sqrt{3} }{ 4 }}=\frac{ \sqrt{2-\sqrt{3}} }{ 2 }\]

OpenStudy (anonymous):

cos (5 pi/12) = sqrt( ( (2-sqrt(3))/4 ) ) Would that be the final answer ?

OpenStudy (misty1212):

yes, but not really because you know that the square root of four is two

OpenStudy (misty1212):

so you would probably write \[\frac{\sqrt{2-\sqrt3}}{2}\]

OpenStudy (anonymous):

Oh so the final answer is 1/2 ?

OpenStudy (misty1212):

oh no !! the final answer is what i wrote above

OpenStudy (anonymous):

Oh lol I thought I had to simplify it

OpenStudy (anonymous):

Thank you so much misty1212 This Really helped me a lot!! :D